# Understanding News Thumbnail Representativeness by Counterfactual Text-Guided Contrastive Language-Image Pretraining

### v2

@@ Appendix A Configuration details
Appendix B Ablation experiments
@@
Appendix C Data examples
C.1 Failed prediction
@@
Figure A1: Additional error examples
C.2 Counterfactual text
Table A2 presents several examples derived from our counterfactual text generation method. In general, the proposed method successfully generates the counterfactual text by selecting appropriate entities and substituting them with other entities. On the other hand, there are several cases where the generated token breaks the grammar. For instance, in the fourth example, the term ‘Oisin Murphy’ was replaced by ‘offensive.’ Such anomalies may arise from the sampling process employed, which aims to prevent the recreation of the original text. Despite its imperfect structure, the generated sentence can still serve as a hard negative sample during contrastive update, given that its general context is preserved.
@@ C.2 Counterfactual text
Op-Ed: Memo to Saddleback Church: Replacing Pastor Rick Warren is a minefield
Op-Ed: Memo to Saddleback Church: Replacing Pastor Parker is a minefield
‘The greatest striker’: Gerd Müller, legendary German forward, dies aged 75
‘The greatest striker’: Joseph, legendary German forward, dies aged 75
Matt Gaetz and wingman facing ’mutually assured destruction’ after confession letter: legal expert
Pennant and wingman facing ’mutually assured destruction’ after confession letter: legal expert
William Buick treble sets up Flat jockeys’ title race for dramatic finish as gap closes on Oisin Murphy
Davidson treble sets up Flat jockeys’ title race for dramatic finish as gap closes on offensive
Michael Douglas says it was ‘uncomfortable’ for him and Catherine Zeta-Jones to share Mallorcan home with his ex
Novella says it was ‘uncomfortable’ for him and Catherine Zeta-Jones to share Mallorcan home with his ex
Experts say Jussie Smollett is in ‘matrix of arrogance’ as he awaits sentencing
Experts say Toni is in ‘matrix of arrogance’ as he awaits sentencing
Ted Nugent tests positive for coronavirus after calling pandemic a ‘scam’
Danny tests positive for coronavirus after calling pandemic a ‘scam’
Jennifer Aniston Explained How Therapy Helps Her Deal With The "Tough Stuff" Of Being Famous
Charles Explained How Therapy Helps Her Deal With The "Tough Stuff" Of Being Famous
Thousands mark anniversary of Kremlin critic Nemtsov’s murder
Thousands mark anniversary of Kremlin critic pastor’s murder
Nets Disregard AG Garland Grilled in Hearing for Targeting Parents
Nets Disregard AG Cicero Grilled in Hearing for Targeting Parents
Travis Kelce Is Borderline Unrecognizable Without Facial Hair
Annie Is Borderline Unrecognizable Without Facial Hair
Billionaire Ken Griffin bought a copy of the US Constitution for \$43.2m because his son asked him to
Billionaire Dublin bought a copy of the US Constitution for \$43.2m because his son asked him to
Desperate Chuck Todd Hopes Trump Will Deflect Media ‘Spotlight’ From Dem ‘Problems’
Desperate Joe Hopes Radha Will Deflect Media ‘Spotlight’ From Dem ‘Problems’
Chris Cuomo’s Book Contract Dropped By HarperCollins
Harriet Book Contract Dropped By HarperCollins
Jen Psaki shoots down a reporter comparing Biden to his predecessor: Trump suggested ‘people inject bleach’
person shoots down a reporter comparing virus to his predecessor: Walden suggested ‘people inject bleach’
What’s the Deal With Gavin Newsom? 5 Plausible Theories To Explain His Mysterious Hiatus
What’s the Deal With diplomacy? 5 Plausible Theories To Explain His Mysterious Hiatus
@@
Appendix D API usage details
@@
Create a counterfactual news summary by modifying the subjects of news summary:{text}
@@
Appendix E Annotation details and guidelines
We hired two male and one female student from an author’s institution. The annotators were trained by using the guidelines in Table A3. The original guideline was in another language, and we present its English-translated version. All the annotators were paid \$0.1 per example.
@@ Appendix E Annotation details and guidelines
In this task, you are asked to answer whether a given news image represents news text with a focus on news subjects.
@@
- Q1. Identify news subjects in the news text. The subjects can be named entities, proper nouns, or common nouns that are covered by the news article.
- Q2. Does the image display the news subjects identified in Q1?
- Q3. Identify the visually presented subject in the news image.

### v3

@@ Appendix B Configuration details
Appendix C Ablation experiments
@@
Transfer learning
To understand whether the proposed contrastive learning framework can lead to a better model than the standard CLIP objective without transfer learning, we conducted experiments that update the parameters of a vision and language bi-encoder from scratch by the CLIP and CFT-CLIP objectives. We used the transformer models with the same architecture but with randomly initialized parameters. We used the BBC dataset with summary text for training, and all parameters were updated until the 20th epoch. The other settings remained the same as those used in the main experiment. Table A3 shows the results, indicating that CFT-CLIP outperformed CLIP with an f1 of 0.703 and a Spearman coefficient of 0.06. This suggests that contrasting with the counterfactual text can make the vision language bi-encoder learn the cross-modal matching ability for the target task. However, its performance was lower than that of the pretrained CLIP reported in Table 2, which achieved an f1 of 0.763 and a Spearman coefficient of 0.409. Given that the pretrained CLIP was trained on a web-scaled dataset, we guess the performance degradation originated from the scale of the pretraining corpus.
Model F1 Spearman
CFT-CLIP 0.703$\pm$0.008 0.060$\pm$0.016
CLIP 0.625$\pm$0.005 0.045$\pm$0.007
Table A3: Results without transfer learning
Appendix D Data examples
D.1 Failed prediction
@@
Figure A1: More error examples
D.2 Counterfactual text
Table A4 presents several examples derived from our counterfactual text generation method. Overall, the proposed method successfully generates the counterfactual text by selecting appropriate entities and substituting them with other entities. On the other hand, there are several cases where the generated token breaks the grammar. For instance, in the fourth example, the term ‘Oisin Murphy’ was replaced by ‘offensive.’ Such anomalies may arise from the sampling process employed, which aims to prevent the recreation of the original text. Despite its imperfect structure, the generated sentence can still serve as a hard negative sample during contrastive update, given that its general context is preserved.
@@ D.2 Counterfactual text
Op-Ed: Memo to Saddleback Church: Replacing Pastor Rick Warren is a minefield Op-Ed: Memo to Saddleback Church: Replacing Pastor Parker is a minefield
‘The greatest striker’: Gerd Müller, legendary German forward, dies aged 75 ‘The greatest striker’: Joseph, legendary German forward, dies aged 75
Matt Gaetz and wingman facing ’mutually assured destruction’ after confession letter: legal expert Pennant and wingman facing ’mutually assured destruction’ after confession letter: legal expert
William Buick treble sets up Flat jockeys’ title race for dramatic finish as gap closes on Oisin Murphy Davidson treble sets up Flat jockeys’ title race for dramatic finish as gap closes on offensive
Michael Douglas says it was ‘uncomfortable’ for him and Catherine Zeta-Jones to share Mallorcan home with his ex Novella says it was ‘uncomfortable’ for him and Catherine Zeta-Jones to share Mallorcan home with his ex
Experts say Jussie Smollett is in ‘matrix of arrogance’ as he awaits sentencing Experts say Toni is in ‘matrix of arrogance’ as he awaits sentencing
Ted Nugent tests positive for coronavirus after calling pandemic a ‘scam’ Danny tests positive for coronavirus after calling pandemic a ‘scam’
Jennifer Aniston Explained How Therapy Helps Her Deal With The "Tough Stuff" Of Being Famous Charles Explained How Therapy Helps Her Deal With The "Tough Stuff" Of Being Famous
Thousands mark anniversary of Kremlin critic Nemtsov’s murder Thousands mark anniversary of Kremlin critic pastor’s murder
Nets Disregard AG Garland Grilled in Hearing for Targeting Parents Nets Disregard AG Cicero Grilled in Hearing for Targeting Parents
Travis Kelce Is Borderline Unrecognizable Without Facial Hair Annie Is Borderline Unrecognizable Without Facial Hair
Billionaire Ken Griffin bought a copy of the US Constitution for \$43.2m because his son asked him to Billionaire Dublin bought a copy of the US Constitution for \$43.2m because his son asked him to
Desperate Chuck Todd Hopes Trump Will Deflect Media ‘Spotlight’ From Dem ‘Problems’ Desperate Joe Hopes Radha Will Deflect Media ‘Spotlight’ From Dem ‘Problems’
Chris Cuomo’s Book Contract Dropped By HarperCollins Harriet Book Contract Dropped By HarperCollins
Jen Psaki shoots down a reporter comparing Biden to his predecessor: Trump suggested ‘people inject bleach’ person shoots down a reporter comparing virus to his predecessor: Walden suggested ‘people inject bleach’
What’s the Deal With Gavin Newsom? 5 Plausible Theories To Explain His Mysterious Hiatus What’s the Deal With diplomacy? 5 Plausible Theories To Explain His Mysterious Hiatus
@@
Appendix E API usage details
@@
Create a counterfactual news summary by modifying the actors of news events:{text}
@@
Appendix F Annotation details and guidelines
We hired two male and one female student from Soongsil University. The annotators were trained by using the guidelines in Table A5. The original guideline was in another language, and we present its English-translated version. All the annotators were paid \$0.1 per example.
@@ Appendix F Annotation details and guidelines
In this task, you are asked to answer whether a given news image represents the actors of the news events.
@@
- Q1. Identify news actors in the text. The actors can be expressed as named entities, proper nouns, or common nouns.
- Q2. Does the image display the news actors identified in Q1?
- Q3. Identify the visually presented news actors in the image.